OpenStudy (perl):

Criticise the following 'proof' that 3=2 x^3 = x*x^2 = x^2 + x^2 + ... + x^2 differentiate both sides 3x^2 = 2x + 2x + 2x + ... + 2x 3x^2 = 2x * x 3x^2 = 2x^2 3 =2

OpenStudy (compassionate):

 If I have 3 apples in one hand, and 2 in the other, then I have more apples in the hand with 3 apples! :)

OpenStudy (perl):

thats true :)

OpenStudy (compassionate):

You have to think; the proof is semantics. It's the way you look at things.

OpenStudy (compassionate):

x*x^2 =/= x^2 + x^2 because that implies 2x^2 = x*x^2

OpenStudy (perl):

i wrote x*x^2 = x^2 + x^2 + ... + x^2 ,

OpenStudy (perl):

there are x terms of x^2

OpenStudy (perl):

comp , you are on the right track though

OpenStudy (compassionate):

Huh. I am all too tired. I bet @dan815 , @myininaya , @whpalmer4 , @zepdrix would love to have a crack at this. Also @wio

OpenStudy (perl):

or @ganeshie8

ganeshie8 (ganeshie8):

I know another flavor of this proof :) x^3 = `x`*x^2 = x^2 + x^2 + ... + x^2 here, number of terms on right hand side depends on x but the proof is treating it as constant. you need to use chain rule for the right hand side while differentiating